A stone falls freely under gravity. It covers distances h1, h2 and h3 in the first 5 seconds, the next 5 seconds and the next 5 seconds respectively. The relation between h1, h2 and h3 is - (1) h1 = 2h2 = 3h3 (2) h1 = 5 h 3 h2 3 = (3) h2 = 3h1 and h3 = 3h2 (4) h1 = h2 = h3 

Grade:11

2 Answers

grenade
8 years ago
(2)
Distance covered in first 5 sec. h1 = 0 + 2 1 a(5)2 h1 = 2 25a …(1) distance covered in first 10 sec S2 = 0 + 2 100a a(10) 2 1 2 = So distance covered in second 5 sec. h2 = S2 – h1 = 2 75a 2 25a 2 100a − = …(2) distance covered in first 15 sec. S3 = 0 + 2 225a a(15) 2 1 2 = so distance covered in last 5 sec. h3 = S3 – S2 = 2 125a 2 100a 2 225a − = ...(3) using (1), (2) and (3) equation. 2 125a h 2 75a h 2 25a h1 2 3 = = h1 = 5 h 3 h2 3 =
 
plz aprve
Rutuja kale
6 years ago
S=ut+1/2gt2 h1=5u+1/2×10×(5)2h1=5u+125h1+h2=10u+500=> h2=5u+375h1+h2+h3=15u+1125=>h3=5u+625Here,u=0h1=125,h2=375,h3=625h1=h2/3=h3/5